OpenStudy (anonymous):

PLEASE HELP! I dont get this at all!!! :( The resistance of a wire varies directly with its length and inversely with the square of its diameter .01 inch as a resistance of 10 ohms, what is the resistance of 60 feet of the same type of wire if its diameter is .02 inch?

jimthompson5910 (jim_thompson5910):

The resistance of a wire varies directly with its length and inversely with the square of its diameter means that \[\Large r = \frac{kL}{d^2}\] where r is the resistance, L is the length, d is the diameter, and k is some fixed number

jimthompson5910 (jim_thompson5910):

Something seems missing though

OpenStudy (anonymous):

in the problem something is missing?

jimthompson5910 (jim_thompson5910):

can you post a screenshot of the problem?

OpenStudy (anonymous):

yesss

jimthompson5910 (jim_thompson5910):

ok thx

OpenStudy (anonymous):

im about to send it, its #9

jimthompson5910 (jim_thompson5910):

ok i see it now, thx

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

it says that 100 ft of wire with a diameter of 0.01 inches has a resistance of 10 ohms so all this means L = 100, d = 0.01, and r = 10 plug all this in and solve for k \[\Large r = \frac{kL}{d^2}\] \[\Large 10 = \frac{k*100}{0.01^2}\] \[\Large 10 = \frac{k*100}{0.0001}\] \[\Large 10 = 1,000,000k\] \[\Large \frac{10}{1,000,000} = k\] \[\Large k = \frac{10}{1,000,000}\] \[\Large k = \frac{1}{100,000}\] \[\Large k = 0.00001\]

jimthompson5910 (jim_thompson5910):

So \[\Large r = \frac{kL}{d^2}\] turns into \[\Large r = \frac{0.00001L}{d^2}\]

jimthompson5910 (jim_thompson5910):

Now plug in L = 60 and d = 0.02 and evaluate to find the final answer

jimthompson5910 (jim_thompson5910):

\[\Large r = \frac{0.00001L}{d^2}\] \[\Large r = \frac{0.00001*60}{0.02^2}\] \[\Large r = ?\]

OpenStudy (anonymous):

how did you get 1,000,000 as a denominator???

jimthompson5910 (jim_thompson5910):

i divided both sides by 1,000,000

jimthompson5910 (jim_thompson5910):

i got 1,000,000 from 100 divided by 0.0001

OpenStudy (anonymous):

ohhhh okay got it

jimthompson5910 (jim_thompson5910):

so what do you get for r

OpenStudy (anonymous):

im doin it

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

i got 15

jimthompson5910 (jim_thompson5910):

that seems off, hold on

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

I'm getting this \[\Large r = \frac{0.00001L}{d^2}\] \[\Large r = \frac{0.00001*60}{0.02^2}\] \[\Large r = \frac{0.0006}{0.0004}\] \[\Large r = 1.5\]

jimthompson5910 (jim_thompson5910):

so you lost a zero somewhere, but other than that, it looks good

OpenStudy (anonymous):

oops i only had 0.006 not 0.0006

jimthompson5910 (jim_thompson5910):

ah that makes more sense

OpenStudy (anonymous):

yeahh, my bad! so is there anything i need to do for the problem???

OpenStudy (anonymous):

yeahh, my bad! so is there anything i need to do for the problem???

OpenStudy (anonymous):

ohhh thats it

jimthompson5910 (jim_thompson5910):

nope, you found the resistance of the 60 ft wire with a diameter of 0.02 inches so you're done

OpenStudy (anonymous):

thank you sooo much !!!

jimthompson5910 (jim_thompson5910):

yw